# Interchangeable Token Embeddings for Extendable Vocabulary and Alpha-Equivalence

### v2

@@ A.2 Temporal logic overview
An LTL formula is evaluated over a trace, which represents a sequence of truth values for atomic propositions over time. In this work, as in DeepLTL [Hahn et al., 2021], we consider symbolic traces of infinite length. These traces are expressed in what is known as a lasso form, denoted $uv^{\omega}$, where $u$ is a finite prefix, and $v$ is a finite sequence that repeats indefinitely.
@@
For compatibility with the dataset from DeepLTL [Hahn et al., 2021], both our traces and formulae are represented in Polish (prefix) notation, where operators precede their operands. For instance, $a\wedge b$ is written as \&ab, which avoids the need for parentheses to resolve ambiguities.
@@
A.3 Hyperparameters
The constant hyperparameter choices for all experiments are given in Table 4. These hyperparameters are kept constant within an experiment. The hyperparameters for the LTL task is taken from DeepLTL [Hahn et al., 2021].
@@ A.3 Hyperparameters
Copy (Sections 5.1.1 and 5.1.2) 64 2 4 64 512 20K
Copy (Section 5.1.4) 128 6 8 128 512 20K
@@
A.3.1 Hyperparameter Search
On the smaller copying task, we train multiple models that use different random embedding methods (Section 4.2) with different $d_{\beta}$ values. While altering $d_{\beta}$, we keep the total number of embedding dimensions $d_{\alpha}+d_{\beta}$ constant. We train each model at least 3 times with different seeds and report the results for the best one in Tables 5 (proposed method) and 6 (baselines).
@@ A.3.1 Hyperparameter Search
The results in Tables 5 and 6 exhibit high variance with no clear patterns that indicate which methods are better. Therefore, we perform an analysis based on correlation coefficients between these hyperparameters and the edit distance using the results from all 277 models we’ve trained (not including the baseline models). For this analysis, we assume that the value of Boolean properties (such as $f_{bn}$, $f_{fn}$ and AdaCos) are 0 or 1. The correlation coefficients are as follows:
@@
A.4 LTL Experiment with Limited Dataset
@@
The result of limiting the number of training samples is similar to the dataset perturbation, albeit much less pronounced for the baseline model. Instead of seeing the performance of the baseline model plummet as in the perturbation experiment, we observe that all models trained on the limited dataset perform similarly in terms of correctness ratio. The biggest difference is observed in the alpha-covariance values, particularly in the 5 AP category, whose ranking aligns with the perturbation experiment.
Table 7: Evaluation of the baselines and our method trained on different versions of LTLRandom35. The same results from Table 2 are shown for easier comparison. The alpha-renaming baseline was trained using 5 AP embeddings since vocabulary generalization is not valuated here. First two columns denote the training dataset and the model. Next two columns indicate the ratio of the correct predictions and exact matches on 99,989 test set samples as evaluated by spot. Last three columns display mean alpha-covariance values for varying atomic proposition (AP) counts, evaluated on all alpha-equivalent variants of 1000 test samples.

### v3

@@ Appendix B Temporal logic overview
An LTL formula is evaluated over a trace, which represents a sequence of truth values for atomic propositions over time. In this work, as in DeepLTL (Hahn et al., 2021), we consider symbolic traces of infinite length. These traces are expressed in what is known as a lasso form, denoted $uv^{\omega}$, where $u$ is a finite prefix, and $v$ is a finite sequence that repeats indefinitely.
@@
For compatibility with the dataset from DeepLTL (Hahn et al., 2021), both our traces and formulae are represented in Polish (prefix) notation, where operators precede their operands. For instance, $a\wedge b$ is written as &ab, which avoids the need for parentheses to resolve ambiguities.
@@
Appendix C Propositional Logic
Unlike LTL (Appendix B), propositional logic does not feature any temporal operators, but we include the derived operators for equivalence ($\leftrightarrow$) and exclusive or ($\oplus$) alongside the basic negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$). This leads to the syntax given in Equation 4, defined over a finite set of atomic propositions $P$ where $p\in P$ an atomic proposition.
$\phi:=\mathbf{T}\mid p\mid\neg\phi\mid\phi_{1}\wedge\phi_{2}\mid\phi_{1}\vee\phi_{2}\mid\phi_{1}\leftrightarrow\phi_{2}\mid\phi_{1}\oplus\phi_{2}$ (4)
In assignment prediction problem for propositional logic, the goal is to determine a Boolean assignment for every atomic proposition $p\in P$ such that the given formula is satisfied. We allow the assignments to be partial, e.g., just as $a=1$, $b=1$ is a valid assignment for the formula $a\vee b$, so is $a=1$, which allows $b$ to take any value.
To encode the assignments for the neural network, an alternating sequence of atomic propositions and values is used. For example, a1b0 represents the assignment $a=1$ and $b=0$. To verify the outputs of the neural network and to generate datasets, pyaiger was used (Vazquez-Chanlatte & Rabe, ).
Appendix D Hyperparameters
The constant hyperparameter choices for all experiments are given in Table 3. These hyperparameters are kept constant within an experiment. The hyperparameters for the logic tasks are taken from DeepLTL (Hahn et al., 2021). For the LTL task, we used the same hyperparameters. On the other hand, for the propositional logic task, we had to make some changes to adapt them to our updated architecture. Firstly, since we utilize weight sharing, we cannot separate the embedding dimensions of encoder and decoder. As a result, instead of having an embedding dimension of 128 for the encoder and 64 for the decoder, we use 128 for both. However, since there are 6 attention heads, we round it up to 132.
@@ Appendix D Hyperparameters
Copy (Sections E.1 and E.2) 64 2 4 64 512 20K
Copy (Section E.5) 128 6 8 128 512 20K
@@
Propositional Logic (Section 5.3) 132 6 6 512 1024 50K
Appendix E Copying Task Experiments
Evaluation method. We generate the predictions using greedy sampling in the copying task. We use the edit distance between the prediction and the ground truth as our evaluation metric. To generate the evaluation datasets (validation and test splits), we create 100 samples for each possible combination of unique character count and string length, starting from a minimum of 3. Consequently, the total evaluation dataset is arranged in a matrix in which the rows represent unique character count in the string and the columns represent the string length. This matrix is upper triangular since the unique character count cannot exceed the string length. For random embeddings, we repeat the evaluation 10 times and report the average. To evaluate up to the string length of 30 in this setup, $10\times 100\times 406=406000$ predictions are required, where $406$ is the number of upper triangular elements in a $28\times 28$ matrix. To minimize the impact of random factors, we train each model three times and report the results only for the best.
E.1 Generalization to larger vocabularies
E.2 Generalization to larger vocabularies and lengths
We create a dataset consisting of 10 million strings whose lengths vary between 5 and 10 with at most 5 unique characters. We evaluate on the same validation set as before, expecting the model to generalize to both longer lengths and larger vocabulary sizes. In the next subsection, we perform a hyperparameter search over random embedding methods, $d_{\beta}$ values, and whether $f_{bn}$, $f_{fn}$, AdaCos are enabled.
E.3 Hyperparameter Search
On the smaller copying task, we train multiple models that use different random embedding methods (Section 4.2) with different $d_{\beta}$ values. While altering $d_{\beta}$, we keep the total number of embedding dimensions $d_{\alpha}+d_{\beta}$ constant. We train each model at least 3 times with different seeds and report the results for the best one in Tables 4 (proposed method) and 5 (baselines).
@@ E.3 Hyperparameter Search
The results in Tables 4 and 5 exhibit high variance with no clear patterns that indicate which methods are better. Therefore, we perform an analysis based on correlation coefficients between these hyperparameters and the edit distance using the results from all 277 models we’ve trained (not including the baseline models). For this analysis, we assume that the value of Boolean properties (such as $f_{bn}$, $f_{fn}$ and AdaCos) are 0 or 1. The correlation coefficients are as follows:
@@
E.4 Sensitivity to randomness in embeddings
To reduce the computational cost of evaluation in other experiments (All LTL experiments and Section E.5), we generate 10 random embeddings, sort them by their cross entropy loss on the evaluated dataset, and use the median one. We find that this serves as a decent proxy for the average performance. Across the validation set evaluations of all 277 models, the percent difference in edit distance between this median method and the real mean is 1.4% on average (meaning that the result from the median method is worse), and 9.1% if we consider the absolute differences.
E.5 Scaling up
Appendix F LTL Experiment with Limited Dataset
@@
The result of limiting the number of training samples is similar to the dataset perturbation, albeit much less pronounced for the baseline model. Unlike in the perturbation experiment, where the baseline model’s performance plummets, all models trained on the reduced dataset maintain similar correctness ratios. The biggest difference is observed in the alpha-covariance values, particularly in the 5 AP category, whose ranking aligns with the perturbation experiment.
Since LTLRandom35 is a synthetic dataset, it exhibits minimal inherent bias, even when the dataset size is limited. Consequently, limiting the dataset size has a smaller effect than introducing perturbations. Furthermore, since the alpha-renaming model was trained using 5 AP embeddings in this experiment, it loses its vocabulary generalization capability unlike our proposed method. Training the alpha-renaming baseline with more APs would require learning a new embedding for each AP, which would reduce its performance.
Table 6: Evaluation of the baselines and our method trained on different versions of LTLRandom35. The same results from Table 1 are shown for easier comparison. The alpha-renaming baseline was trained using 5 AP embeddings since vocabulary generalization is not evaluated here. First two columns denote the training dataset and the model. Next two columns indicate the ratio of the correct predictions and exact matches on 99,989 test set samples as evaluated by spot. Last three columns display mean alpha-covariance values for varying atomic proposition (AP) counts, evaluated on all alpha-equivalent variants of 1000 test samples.
@@
Appendix G Propositional Logic Experiments
This section gives more details about the experimental setup of the propositional logic task and continues the experiments.
G.1 Experimental Setup
We use PropRandom35 from DeepLTL (Hahn et al., 2021) as our main 5 AP dataset, and create other datasets using the same approach. In particular, propositional logic formulae are generated randomly, with negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$) operators having an equal weight. Equivalence ($\leftrightarrow$) and exclusive or ($\oplus$) operators each have half as much weight since they are derived operators. The corresponding assignment is generated by querying the pyaiger’s SAT solver for a minimal unsatisfiable core of the negated formula.
As in the LTL experiments, we use a transformer encoder-decoder architecture with three-way weight tying (Press & Wolf, 2016). The positional encoding method is tree-positional encoding (Shiv & Quirk, 2019) for the encoder and RoPE (Su et al., 2024) for the decoder. Predictions are generated using beam search with a beam size of 3.
Since the network outputs the assignments as a sequence (Appendix C), the same assignment can be encoded in multiple ways by changing the order. For example, both a1b0 and b0a1 represent the same set of assignments $a=1$ and $b=0$, which can be written as $\{(a,1),(b,0)\}$ in set notation. We consider such pairs exact matches in the propositional logic experiments. If the predicted assignment does not exactly match the ground truth, we use pyaiger to evaluate the correctness.
G.2 Dataset Perturbations
In this section, we repeat the dataset perturbation experiment (Section 5.2.1) for the propositional logic task. The perturbation is introduced in a similar manner by renaming the APs such that the order of the first AP appearances in the label (sequence denoting the Boolean assignment) is always the same. As shown in Table 7, the experimental results once again confirm that our method introduces a robust inductive bias for alpha-equivalence.
Table 7: Evaluation of the baselines, our method, and Llama 3.2 on the PropRandom35 dataset. The alpha-renaming baseline was trained using 5 AP embeddings since vocabulary generalization is not evaluated here. First two columns denote the training dataset and the model. Next two columns indicate the ratio of the correct predictions and exact matches on 100,000 test set samples as evaluated by pyaiger. Last three columns display mean alpha-covariance values for varying atomic proposition (AP) counts, evaluated on all alpha-equivalent variants of 1000 test samples.
Training Evaluation Alpha-Covariance
Dataset Model Correct Exact 3 AP 4 AP 5 AP
Normal Baseline 95.62% 57.94% 95.70% 93.69% 76.02%
Perturbed Baseline 41.57% 9.04% 14.96% 16.85% 10.65%
Perturbed Alpha-Renaming 93.85% 57.24% 99.56% 99.60% 93.23%
Perturbed Proposed 93.25% 56.45% 99.23% 99.42% 92.98%
Pretrained Llama 3.2 3B 29.03% 1.56% 50.75% 27.96% 11.25%
Appendix H Computational Efficiency Details
To evaluate the practical applicability of our method, we analyze its computational overhead compared to baseline approaches. We report training times, inference speeds, and memory requirements across different experimental settings.
Training efficiency. We measured training durations for models trained on NVIDIA H100 GPUs using identical hyperparameter settings. In LTL solving task, the average training times were as follows:
  • Baseline (traditional embeddings): 2 hours 12 minutes
  • Alpha-renaming baseline: 2 hours 33 minutes
  • Proposed method: 2 hours 29 minutes
The proposed method introduces minimal training overhead compared to the baseline, with only a 13% increase in training time. This modest overhead stems from the additional embedding preparation steps required during training.
Inference performance. We conducted a runtime analysis using our best-performing LTL model on NVIDIA A4000 hardware. The model uses Hypercube Vertices randomization with uniqueness checking enabled, evaluated with batch size 768 and beam search (beam size = 3). In this setup, a forward pass takes 0.206 seconds, and autoregressive generation 9.808 seconds. On the other hand, the embedding preparation time is measured at 0.0003 seconds, which is negligible compared to model execution. Importantly, during inference, embeddings need only be generated once at the start of the evaluation session, making the amortized cost even smaller for batch processing.
Memory overhead. Our method reduces the total parameter count compared to traditional approaches since only one common embedding is learned for all interchangeable tokens, regardless of their quantity. The memory overhead comes primarily from constructing the embedding matrix during runtime, which requires temporary storage for the randomized components. However, this additional memory requirement is on the same order of magnitude as the embedding matrix itself, which represents a small fraction of total model parameters in transformer architectures.
The parameter efficiency of our method scales favorably with vocabulary size. Unlike traditional approaches that require learning separate embeddings for each token (thereby scaling linearly with the vocabulary size), our method’s parameter count remains constant regardless of the number of interchangeable tokens. However, two factors require consideration for very large vocabularies:
  1. 1.
    Sampling set size: In discrete random generation methods, the sampling set is naturally bounded (Table 1). However, the sampling set grows exponentially with the number of dimensions, ensuring sufficient diversity even for large vocabularies.
  2. 2.
    Uniqueness checking: For vocabularies with hundreds of thousands of tokens, uniqueness verification becomes computationally expensive, but the probability of collisions decreases exponentially with increasing embedding dimensions.
Appendix I LLM Setup
We use the 3B-parameter version of Llama 3.2 (Grattafiori et al., 2024), quantized with Q4_K_M, and run it using Ollama 0.4.7 as our LLM backend. We first experimented with greedy sampling (by setting top-k=1) since Ollama does not support beam search. However, we found that the default sampling options (top-k=40 and top-p=0.9) yielded better results. Therefore, we use these default settings for all experiments.
Unlike our specialized models, which operate on prefix (Polish) notation, we prompt the LLM using infix notation for input formulas (and output traces in the LTL task), as this format is more prevalent in natural language and more familiar to general-purpose LLMs. To output the assignments in the propositional logic task, we use JSON format, and constrain the LLM’s output using a JSON schema. The input prompts for the LTL and propositional logic tasks are given in Listing LABEL:lst:llm-prompt-ltl and Listing LABEL:lst:llm-prompt-prop, respectively. For each sample, the “{formula}” substring in the prompt is replaced by the input formula, and the prompt is given as a user message to the LLM.
We set the random seed to 42 for each sample. Although the reason behind this choice is reproducability, it also seems to improve alpha-covariance. For example, the alpha-covariance values reported for Llama 3.2 in Table 1 are 68.17%, 63.27%, 62.34% for 3 to 5 APs, respectively, which decrease to 41.94%, 43.10%, 44.62% when the random seed is no longer fixed.
Listing 1: LLM Prompt for the LTL solving task.
1Your task is to generate a satisfying trace for a given LTL (Linear Temporal Logic) formula.
2Lowercase letters denote the atomic propositions.
3The output trace should be in lasso form composed of two parts: the prefix part and the cycle part.
4Timesteps in the trace should be separated by semicolons, and the cycle part should be enclosed in curly braces, preceeded by the keyword "cycle".
5
6Temporal operators:
7X: Next operator
8U: Until operator
9
10Logical operators:
11&: AND operator
12|: OR operator
13!: NOT operator
14The output trace is a symbolic trace, which means that the logical operators are allowed, but not temporal operators.
15
16Constants:
170: False
181: True
19Note that other numbers are invalid.
20
21Example 1
22Formula: X((a & Xa) U XXb)
23Trace: 1; 1; 1; b; cycle{{1}}
24
25Example 2
26Formula: !c U X(1 U b)
27Trace: 1; b; cycle{{1}}
28
29Example 3
30Formula: X!X!(b & Xb)
31Trace: 1; 1; b; b; cycle{{1}}
32
33Example 4
34Formula: !(1 U !c)
35Trace: cycle{{c}}
36
37Your Turn
38Formula: {formula}
39Please generate the corresponding trace. Output the trace only.
Listing 2: LLM Prompt for the propositional logic task.
1Your task is to generate an assignment that satisfies a given propositional logic formula.
2Lowercase letters denote the atomic propositions.
3The output is a JSON object representing the assignment.
4
5Logical operators (ordered from highest precedence to lowest):
6!: NOT operator
7&: AND operator
8|: OR operator
9xor: Exclusive OR operator
10<->: Logical equivalence operator (biconditional)
11
12Constants:
130: False
141: True
15Note that other numbers are invalid.
16
17Example 1
18Formula: !a | c & (b <-> c)
19Assignment: { "a": false }
20
21Example 2
22Formula: !(a <-> (!a xor !e))
23Assignment: { "a": true, "e": true }
24
25Example 3
26Formula: a & (!a <-> !c | d)
27Assignment: { "a": true, "c": true, "d": false }
28
29Example 4
30Formula: !(a | !(!d | b & d))
31Assignment: { "a": false, "d": false }
32
33Your Turn
34Formula: {formula}
35Please generate an assignment that satisfies this formula. Output the assignment only, in JSON format.